OpenStudy (anonymous):

Please Help!! Sequence and Series: Show that the sequence 8, 4√2, 4, 2√2, ... is geometric. Hence find, in simplest form, the general term Un.

OpenStudy (anonymous):

What do you know of geometric series? :)

OpenStudy (anonymous):

i know hw to do it; i just cant simplify it enough to get the right answer. so far i have: u1=8 and r = 1/√2 so the equation i have right now is Un = 8(1/√2)^n-1 I don't know how to simplify it further. The real answer is completely different.

OpenStudy (anonymous):

Is the answer \[\huge u_n=8\cdot2^{\frac{1-n}2} \]?

OpenStudy (anonymous):

no it is \[u_{n}=2^{\frac{ 7 }{ 2 } - \frac{ n }{ 2 }}\]

OpenStudy (anonymous):

The book also says that my values for u1 and r are correct then i dont know what is wrong with my equation

OpenStudy (anonymous):

Well, of course :) \[\huge u_n=8\cdot2^{\frac{1-n}2}=2^3\cdot 2^{\frac{1-n}2}=2^{3+\frac{1-n}{2}}\]

OpenStudy (anonymous):

o i see... omg. Thank you so much @PeterPan :)